An attorney at a major law firm decides to leave his $100,000 per year job and start up his own practice. He invests $50,000 of his own money in his new practice; this money was withdrawn from a bank account that pays an interest rate of 2% per year. The money is used to buy furniture, equipment, the first year’s rent, etc. By the end of his first year, the lawyer determines that his total revenues were $150,000.

a) What is the lawyer’s accounting profit during the first year?

b) What is the lawyer’s economic profit during the first year?

c) Was starting up his own practice the right decision?
Answer #1

Income = 100000

Interest = 2%

money after two year =100000 + (100000*1*2)/100 = 102000

He invests = 50,000

interest on this amount (opportunity cost) = (50000*1*2)/100 = 1000

By the end of first year, his revenue = 150000

a) Accounting profit = revenue - investment

= 150000 - 50000

= 100000

b) Economic profit = revenue - (investment + opprotunity cost)

= 1500000 - 51000

= 99000

c) If he had not given up his job, would have earned= 100000

Now, after starting his own business his profit = 99000

So, starting his own practice was not the right decision.